OpenStudy (dominantvampire):

a person walks 25 degrees south of west for 2 km. how far would he walk due to south? due to west?

OpenStudy (anonymous):

 \[\color{green}{\vec{d}_y = \vec{d} \cdot \cos 65^\circ\\ \vec{d}_x = \vec{d} \cdot \sin 65^\circ}\] The directions will be as shown in the diagram. Use 2 km for vector d when you solve the equations shown above.

OpenStudy (dominantvampire):

wont that b -d

OpenStudy (dominantvampire):

@ghuczek

OpenStudy (dominantvampire):

-d.sin65

OpenStudy (anonymous):

Yes, if you assign negative values for down and left. Exactly.

OpenStudy (dominantvampire):

can u help me out in 1 more

OpenStudy (dominantvampire):

vector B has the magnitude of 20 cm and is directed 25 anti clockwise. from - xaxis express in unit vctr notation

OpenStudy (anonymous):

\[\color{green}{\vec{B}_x = - \vec{B} \cdot \cos 25^\circ\\ \vec{B}_y =- \vec{B} \cdot \sin 25^\circ}\]